The Larger of 2 integers is 4 times the smaller. The sum of the 2 integers is less than twenty. Find the largest possible values for the integers
Answer by Maths68(1474)

Let x be the smaller and y be the larger integer.

The Larger of 2 integers is 4 times the smaller.
y=4x..........(1)
The sum of the 2 integers is less than twenty
x+y<20
Put the value of y from (1) to above equation
x+4x<20
5x<20
5x/5<20/5
x<4
so x is 3,2 or 1
Put the these values of x in (1)
y=4(3)
y=12
y=4(2)
y=8
y=4(1)
y=4
So when x = 3,2,1
y = 12,8,4
Largest possible values (x,y)=(3,12)